# Enhancement of first-order phase transitions through a mass-acquiring scalar field

### v1

@@ Abstract
Phase transitions in the early Universe give rise to effective masses for massless fields in the symmetry-broken phase. We use the lattice simulations to investigate the impact of a spectator scalar field with mass generation on the dynamics of first-order phase transitions and the generation of gravitational waves. In addition to the well-known friction effects, we identify a novel effect that significantly enhances the strength of first-order phase transitions. The amplitude of the mass-acquiring field is highly suppressed in the true vacuum bubbles, resulting in additional release of vacuum energy that concentrate on the bubble walls. We also establish an analytical method to explain our numerical results.
@@ I Introduction
First-order phase transitions (FOPTs) in the early Universe are initiated by the nucleation of true vacuum bubbles through quantum tunneling [1, 2, 3, 4, 5].The subsequent release of vacuum energy accelerates the bubble walls and result in the generation of gravitational waves (GWs) through bubble collisions, sound waves, and turbulence [6, 7, 8, 9, 10, 11, 4, 12]. These transitions represent one of the most promising cosmological origins of stochastic GW backgrounds, potentially observable by upcoming space-based interferometers including LISA [13, 14, 15, 16], Taiji [17, 18], and TianQin [19, 20].
@@
Accurate predictions of the GW energy spectrum require detailed modeling of the underlying dynamics, including the bubble wall acceleration, the energy transfer processes, and the post-collision field evolution. Multiple analytical and numerical techniques have been developed, including the envelope approximation [7, 21], hydrodynamic treatments [9, 22, 23], and fully field-theoretic lattice simulations [24, 25, 26, 27, 28]. These studies have systematically characterized how key physical parameters, particularly the wall thickness, the wall velocity, and the out-of-equilibrium scalar field evolution, determine the spectral features of the resulting GW signals.
@@
More recently, increasing attention has turned to the scenarios where the scalar field that triggers the FOPT couples to additional fields which acquire effective masses in the true vacuum. Such mass-generating mechanism has been incorporated into a variety of beyond Standard Model scenarios, including filtered dark matter [29, 30, 31, 32], Fermi-ball dark matter [33, 34, 35], non-thermal particle production in the dark sector [36, 37], and the Q-ball formation [38, 39]. The presence of mass-acquiring fields also alters the effective potential of FOPTs and lead to energy redistribution. In such cases, even if the mass-acquiring fields are energetically subdominant, they may alter the FOPT dynamics and modify the GW signatures [40].
@@
In this work, we numerically investigate the impact of a mass-acquiring scalar field during an FOPT, focusing on its affect on the bubble wall evolution and the resulting GW energy spectrum. We fix the effective potential for the first-order phase transition (FOPT) that determines the bubble nucleation rate, and numerically compare results across different values of both the coupling constant and the mass-acquiring field fraction. By analyzing simulation results, it was found that the amplitude of the mass-acquiring field within the true vacuum bubbles significantly decreases, which consequently reducing the minimum of the true vacuum potential. This effect is equivalent to an effective enhancement of the transition strength, leading to an increase in bubble wall velocity and an amplification of GWs. We also analyze the redistribution of the released vacuum energy among different components and develop an analytical power-law expression that explains the observed GW amplitude.
@@ II.1 The Effective Potential
On the other hand, the existence of $\chi$ also alter the form of the FOPT potential and change the prediction of the bubble nucleation rate. Since the wavelength of $\chi$ is in general much smaller than the vacuum bubbles, we can apply the mean-field approximation to obtain the reduced effective potential
@@
where $M_{\phi}^{2}\equiv\mu_{\phi}^{2}+2\lambda_{\phi\chi}\langle\delta\chi^{2}\rangle$ is the effective mass of $\phi$ and the angle brackets presents the spatial average. We use the symbol $\delta\chi$ since $\langle\chi\rangle$ the vanishes. Note that Eq. (5) is only responsible for the bubble nucleation, while the numerical calculations of field evolutions are governed by Eq. (3).
For a fast FOPT, the energy scale of the Universe remains roughly unchanged, so quantum tunneling only occurs when $V_{\mathrm{re}}$, or equivalently $M_{\phi}$, reaches the threshold. The parameter $\mu_{\phi}$ generally decreases as the temperature drops. Therefore, the presence of $2\lambda_{\phi\chi}\langle\delta\chi^{2}\rangle$ does not change the effective potential $V_{\mathrm{re}}$ at the the onset of the FOPT, it is the threshold that determines the beginning of sufficient bubble nucleation. For larger $2\lambda_{\phi\chi}\langle\delta\chi^{2}\rangle$, the FOPT is postponed slightly to render a smaller $\mu_{\phi}$, leaving $M_{\phi}$ almost unchanged, as the threshold determines. Consequently, it is reasonable to fix $M_{\phi}$ as a constant among different conditions of $\lambda_{\phi\chi}$ and the amplitude of $\chi$. Furthermore, the direct influence of the interaction terms on the tunneling rate can also be excluded.
@@ III Numerical Methods
We obtain the critical bubble profile using FindBounce, a package that numerically solves the bounce equations via the shooting method [49]. The resulting profile is well fitted by the form of Eq. (7), which corresponds to the analytical solution in the thin-wall approximation. The parameters are set to be $V_{0}=0.2\,\omega_{*}^{2}\eta^{2}$, $\beta=1.3\,\omega_{*}^{2}\eta^{-1}$, and $M_{\phi}^{2}=0.3\,\omega_{*}^{2}$. The corresponding critical bubble radius and wall thickness are $R_{c}=10.06\,\omega_{*}^{-1}$ and $l_{0}=3.05\,\omega_{*}^{-1}$, respectively. At the initial time $t_{i}=0$, $N_{b}$ critical bubbles are nucleated to trigger the FOPT. The centers of any two bubbles are separated by more than $2R_{c}$ to avoid initial overlap. However, Equation (7) indicates that the bubble profile exhibits a tail extending to infinity. To self-consistently place multiple bubbles without introducing sharp cutoffs, the bubbles are initialized by superimposing the profiles as [24]
@@ VI Summary and Discussion
We also establish an analytical formulism that accurately explains our results of the GW energy spectrum. These findings underscore that coupled fields with dynamically evolving mass can significantly backreact to the vacuum structure and field dynamics. This necessitates their inclusion in theoretical predictions of GW signals from early-Universe FOPTs.

### v2

# Dynamical backreaction of a mass-acquiring scalar field on first-order phase transitions

@@ Abstract
Phase transitions in the early Universe give rise to effective masses for massless fields in the symmetry-broken phase. We perform lattice simulations to study the dynamical impact of a mass-acquiring spectator field on the evolution of first-order phase transitions and the associated gravitational-wave production, while keeping the effective potential responsible for bubble nucleation fixed. In addition to the well-known friction effects, we identify a novel effect that significantly enhances the strength of first-order phase transitions. In contrast to the general scenario, although the effective potential governs the tunneling rate, the amplitude of the $\chi$ field is strongly suppressed inside the true vacuum bubble, resulting in a faster bubble expansion than predicted by the effective potential alone. The amplitude of the mass-acquiring field is highly suppressed in the true vacuum bubbles, resulting in additional release of vacuum energy that concentrate on the bubble walls. We further develop an analytical framework that not only explains our numerical results but can also be used to improve the estimation of gravitational-wave signals in related phase-transition scenarios.
@@ I Introduction
First-order phase transitions (FOPTs) in the early Universe are initiated by the nucleation of true vacuum bubbles through quantum tunneling [1, 2, 3, 4, 5, 6, 7]. The subsequent release of vacuum energy accelerates the bubble walls and result in the generation of gravitational waves (GWs) through bubble collisions, sound waves, and turbulence [8, 9, 10, 11, 12, 13, 4, 14]. These transitions represent one of the most promising cosmological origins of stochastic GW backgrounds, potentially observable by upcoming space-based interferometers including LISA [15, 16, 17, 18], Taiji [19, 20], and TianQin [21, 22].
@@
Accurate predictions of the GW energy spectrum require detailed modeling of the underlying dynamics, including the bubble wall acceleration, the energy transfer processes, and the post-collision field evolution. Multiple analytical and numerical techniques have been developed, including the envelope approximation [9, 23], hydrodynamic treatments [11, 24, 25], and fully field-theoretic lattice simulations [26, 27, 28, 29, 30, 31, 32]. These studies have systematically characterized how key physical parameters, particularly the wall thickness, the wall velocity, and the out-of-equilibrium scalar field evolution, determine the spectral features of the resulting GW signals.
@@
More recently, increasing attention has turned to the scenarios where the scalar field that triggers the FOPT couples to additional fields which acquire effective masses in the true vacuum. Such mass-generating mechanism has been incorporated into a variety of beyond Standard Model scenarios, including filtered dark matter [33, 34, 35, 36], Fermi-ball dark matter [37, 38, 39], non-thermal particle production in the dark sector [40, 41], highly interactive particle relics dark matter [42], and the Q-ball formation [43, 44]. The presence of mass-acquiring fields also alters the effective potential of FOPTs and lead to energy redistribution. In such cases, even if the mass-acquiring fields are energetically subdominant, they may alter the FOPT dynamics and modify the GW signatures [45].
In this work, we numerically investigate the impact of a mass-acquiring scalar field during an FOPT, focusing on its effect on the bubble wall evolution and the resulting GW energy spectrum. Throughout this paper the phase transition is driven solely by the scalar field $\phi$. Both the false and true vacua satisfy $\langle\chi\rangle=0$, so the transition proceeds as
$(\phi,\chi):(0,0)\rightarrow(\phi_{\mathrm{b}},0),$
and the $O(4)$-symmetric bounce trajectory is confined to the $\chi=0$ axis. Our setup therefore does not realise a two-field phase transition with non-trivial tunnelling in the $(\phi,\chi)$ plane; instead, $\chi$ plays the role of a spectator field.
@@
We fix the effective potential for the first-order phase transition (FOPT) that determines the bubble nucleation rate, and numerically compare results across different values of both the coupling constant and the mass-acquiring field fraction. This setup allows us to isolate the dynamical effects induced by the field evolution from the trivial modifications of the potential shape, ensuring that the observed differences do not originate from a change in the effective potential itself. By analyzing simulation results, it was found that the amplitude of the mass-acquiring field within the true vacuum bubbles significantly decreases, which consequently reducing the minimum of the true vacuum potential. This effect is equivalent to an effective enhancement of the transition strength, leading to an increase in bubble wall velocity and an amplification of GWs. We also analyze the redistribution of the released vacuum energy among different components and develop an analytical power-law expression that explains the observed GW amplitude.
@@ II.1 The Effective Potential
Because the interaction term does not induce a vacuum expectation value for $\chi$, the relevant minima of $V(\phi,\chi)$ for the phase transition are located at $(\phi,\chi)=(0,0)$ and $(\phi_{\mathrm{b}},0)$, with $\langle\chi\rangle=0$ in both phases. Any $O(4)$-symmetric bounce solution interpolating between these vacua therefore lies entirely on the $\chi=0$ axis. Solving the tunnelling problem with the full two-field potential $V(\phi,\chi)$ is then equivalent to using the one-dimensional potential $V(\phi)$: the field $\chi$ is effectively kept frozen, the Euclidean action and the critical bubble profile coincide with the standard single-field result, and the effect of $\chi$ fluctuations on the tunnelling action and on the nucleation conditions is not captured. For this reason we do not use $V(\phi,\chi)$ or $V(\phi)$ to compute the bounce, but instead adopt the reduced potential $V_{\mathrm{re}}(\phi)$ introduced below, in which the leading impact of $\chi$ fluctuations enters through $\langle\delta\chi^{2}\rangle$.
@@
On the other hand, the existence of $\chi$ also alters the form of the FOPT potential and changes the prediction of the bubble nucleation rate. Since the wavelength of $\chi$ is in general much smaller than the vacuum bubbles, we can apply the mean-field approximation to obtain the reduced effective potential
@@
where $M_{\phi}^{2}\equiv\mu_{\phi}^{2}+2\lambda_{\phi\chi}\langle\delta\chi^{2}\rangle$ is the effective mass of $\phi$ and the angle brackets presents the spatial average. We use the symbol $\delta\chi$ since $\langle\chi\rangle$ vanishes. Note that Eq. (5) is only responsible for the bubble nucleation, while the numerical calculations of field evolutions are governed by Eq. (3). It is worth noting that this analysis is applicable to the weak-coupling regime, where the interaction strength $\lambda_{\phi\chi}$ is small. In this limit, loop effects are subdominant and can be safely neglected within the perturbative regime considered in this work [46, 1, 47].
For a fast FOPT, the energy scale of the Universe remains roughly unchanged, so quantum tunneling only occurs when $V_{\mathrm{re}}$, or equivalently $M_{\phi}$, reaches the threshold. The parameter $\mu_{\phi}$ generally decreases as the temperature drops. Therefore, the presence of $2\lambda_{\phi\chi}\langle\delta\chi^{2}\rangle$ does not change the effective potential $V_{\mathrm{re}}$ at the the onset of the FOPT, it is the threshold that determines the beginning of sufficient bubble nucleation. For larger $2\lambda_{\phi\chi}\langle\delta\chi^{2}\rangle$, the FOPT is postponed slightly to render a smaller $\mu_{\phi}$, leaving $M_{\phi}$ almost unchanged, as the threshold determines. Consequently, it is reasonable to fix $M_{\phi}$ as a constant among different conditions of $\lambda_{\phi\chi}$ and the amplitude of $\chi$. Furthermore, the direct influence of the interaction terms on the tunneling rate can also be excluded. In other words, the effective potential responsible for bubble nucleation is kept unchanged when introducing the additional scalar field $\chi$, so that the subsequent differences in the evolution of fields and the GW signal originate purely from the dynamical evolution of the coupled fields rather than from a static modification of the potential. As we will see, as the evolution proceeds, the expansion of the bubble wall is not dictated by $V_{\mathrm{re}}(\phi)$, but instead aligns more closely with the prediction of $V(\phi)$.
@@ III Numerical Methods
We obtain the critical bubble profile using FindBounce, a package that numerically solves the bounce equations in the reduced potential $V_{\mathrm{re}}(\phi)$ via the shooting method [56]. The resulting profile is well fitted by the form of Eq. (7), which corresponds to the analytical solution in the thin-wall approximation. The parameters are set to be $V_{0}=0.2\,\omega_{*}^{2}\eta^{2}$, $\beta=1.3\,\omega_{*}^{2}\eta^{-1}$, and $M_{\phi}^{2}=0.3\,\omega_{*}^{2}$. The corresponding critical bubble radius and wall thickness are $R_{c}=10.06\,\omega_{*}^{-1}$ and $l_{0}=3.05\,\omega_{*}^{-1}$, respectively. At the initial time $t_{i}=0$, $N_{b}$ critical bubbles are nucleated to trigger the FOPT. The centers of any two bubbles are separated by more than $2R_{c}$ to avoid initial overlap. However, Equation (7) indicates that the bubble profile exhibits a tail extending to infinity. To self-consistently place multiple bubbles without introducing sharp cutoffs, the bubbles are initialized by superimposing the profiles as [27]
@@ VI Summary and Discussion
We also establish an analytical formulism that accurately explains our results of the GW energy spectrum. These findings underscore that coupled fields with dynamically evolving mass can significantly backreact to the vacuum structure and field dynamics. This necessitates their inclusion in theoretical predictions of GW signals from early-Universe FOPTs. This mechanism is generally relevant for scenarios involving mass-acquiring spectator fields, such as filtered dark matter and Q-ball formation. In such cases, analytical estimates of the gravitational-wave spectrum that rely solely on the initial potential may underestimate the signal strength. The lattice approach developed here provides a consistent framework to capture both the trivial potential modification and the additional dynamical energy transfer during the phase transition.
For scenarios involving larger couplings, one-loop quantum corrections to the effective potential may become non-negligible. These corrections introduce additional terms proportional to $\phi^{4}$, $\phi^{2}\chi^{2}$, and $\chi^{4}$ [46, 47]. The first two can be absorbed into the renormalized potential, while the $\chi^{4}$ term may affect the evolution of $\chi$ only when the coupling is strong and its fluctuations are large. A systematic investigation of these effects in the strong-coupling regime is left for future work.
@@
ACKNOWLEDGMENTS
This work is supported in part by the National Natural Science Foundation of China No. 12235019 and No. 12475067, in part by the National Natural Science Foundation of China No. 12147103, No. 12235019, No. 12075297 and No. 12347103.